# Luminous Fast Blue Optical Transients as very massive star core-collapse events

### v1

@@ 4.1 H and He-poor spectra
LFBOTs display H and He poor spectra at early times (e.g. Prentice et al., 2018). The progenitor star, or star being tidally disrupted, therefore must have a He or He-poor composition. In the TDE scenario, a white dwarf is favoured (Kuin et al., 2019). For a BH-stellar merger, the stellar object must be a Wolf-Rayet star (Metzger, 2022). In the scenario explored in this paper, the median H and He mass fractions in the ‘leftover mass’111I.e. the stellar mass leftover after remnant formation. In a successful supernova with no fallback, this would be the ejecta mass. of the models selected as progenitors of BHs with M⊙ are X0.003 and Y0.08 respectively (where the uncertainties are defined by the 16th and 84th percentiles). The distributions of Xsurf and Ysurf are shown in Figure 3. 69% of our selected models are classified as Wolf-Rayet stars given the Xsurf values and high surface temperatures of log10(T/K)4.45 (Eldridge et al., 2017). The outer layers of these stars therefore meet the criteria for being H and He poor, and successfully exploding stars with such low X and Y surface mass fractions are expected to produce type Ic supernovae (i.e. SNe without detectable H and He lines, e.g. Dessart et al., 2012; Eldridge et al., 2013; Chrimes et al., 2020).
@@ 4.2 Dense circumstellar media and radio emission
A key characteristic of LFBOTs is their slow-rising, luminous radio emission, attributed to self-absorbed synchrotron emission from a blast-wave expanding through the circumstellar medium (CSM). These blast-waves are mildly relativistic (e.g. Coppejans et al., 2020), and probe the CSM at radii of 1016–1017 cm. LFBOT radio emission is well described by blast-wave propagation through dense CSM profiles defined by values of 0.1-10, in units of M⊙ yr-1/1000 km s-1. For Wolf-Rayet wind speeds of order 1000 km s-1, radii of 1016–1017 cm corresponds to mass loss in the final decades before explosion. We calculate the mean final values for the selected models, noting that bpass only runs until the end of core carbon burning. Mass loss rates are provided as a bpass output. We adopt standard bpass wind speed prescriptions, namely those of Vink et al. (2001), and Nugis & Lamers (2000) for Wolf-Rayets (we refer the reader to Eldridge et al., 2017; Chrimes et al., 2022, for more details). We find a median final = (0.04) M⊙ yr-1/1000 km s-1. The full distribution of is shown in Figure 3. These mass-loss rate and wind speed values are consistent with the lower end of the values inferred from LFBOT radio observations, which span 0.1–10 (e.g. Yao et al., 2022). We note that mass loss in the final decades before core-collapse may be significant and eruptive, plausibly explaining the non- profiles observed (e.g. Smith, 2014). However, because the stellar evolution models used end after core carbon burning, a full modelling of the CSM around these stars on the relevant scales is beyond the scope of this paper.
@@ 4.3 Long-lived ultraviolet plateaus
Finally, we ask whether the selected stellar models can produce a long-lived tail of UV emission lasting at least several years after the event, as observed in AT2108cow (Sun et al., 2022, 2023; Chen et al., 2023; Inkenhaag et al., 2023). This has been interpreted as due to a long-lived accretion disc, reminiscent of the emission seen in TDEs (Inkenhaag et al., 2023; Mummery & van Velzen, 2025). To see whether our models can produce AT 2018cow-like UV plateaus through this mechanism, we first examine the remaining mass in the selected bpass models which survives outside the nascent BHs (i.e., the pre-core-collapse mass minus the initial BH mass). In Figure 4 we show 2D histograms of these masses at each metallicity, for the fiducial Langer & Norman (2006) CSFH. For more details on the BH mass distribution in bpass, we refer the reader to Eldridge et al. (2017), Ghodla et al. (2022) and Briel et al. (2023).
@@
As outlined in Section 2, bpass calculates BH masses by injecting a standard supernova energy of 1051 erg to determine the mass of the material which is unbound. The remainder is the remnant mass. In a failed supernova, the ‘leftover’ material is not successfully ejected. It may therefore be reasonable to assume that all of the leftover material falls back and is accreted instead, in line with other prescriptions for final BH masses in the massive to very massive star regime, which have close to a one-to-one relation between the pre-core-collapse mass and final remnant mass (after fallback) for stripped envelope progenitors (Fryer et al., 2012; Briel et al., 2023).
With the BH mass, leftover mass, and angular momentum of the leftover mass at the radius of the pre-collapse stellar surface as inputs, we use the model of Mummery & Balbus (2020); Mummery et al. (2024) to calculate UV light-curves for every selected stellar model. The results for the F225W Wide Field Camera 3 (WFC3) Hubble Space Telescope UV filter are shown in Figure 6. We adopt two example scenarios. In the first scenario, all of the leftover material is accreted (as motivated above). This is the ‘100%’ case. We also explore a contrasting scenario in which only 10% of the leftover material is accreted (i.e. 90% is ejected, the 10% case). Massive stars in low metallicity environments such as the Magellanic Clouds are observed to be rotating at 10–20% of their critical velocity (e.g. Ramírez-Agudelo et al., 2013, 2015). In both scenarios, since stellar rotation is not explicitly tracked in bpass, we therefore assume that the star is rotating at 10% of the critical (break-up) velocity. The resulting light-curves primarily lie below the observed late-time F225W points for AT 2018cow, but with a substantial spread, and many models can match the luminosity of AT 2018cow (see also Omand et al. in prep).
@@
We acknowledge that the failed SN interpretation of LFBOTs is predicated on the notion that some non-negligible fraction of the total stellar mass does not participate in the prompt BH formation and instead falls back onto the nascent BH. If the core-collapse of such massive stars actually results in the entire stellar mass being promptly engulfed by the event horizon, there will no material to accrete, ruling out failed supernovae as a viable channel for LFBOTs. However, in our fiducial models this is not the case (see Figure 4).
@@ 5.1 LFBOT environments
In the failed SN scenario, we would expect LFBOTs to be strongly associated with star-forming regions, as the short-lived massive progenitors are not expected to travel far from their birth sites. Although the LFBOT sample is small, it appears as though LFBOTs occupy diverse environments within their hosts, with at least two events at moderate to large host-normalised offsets (Chrimes et al., 2024b; Pursiainen et al., 2025). The apparent preference for LFBOTs to occur in low metallicity environments may help in this regard: in general, galactic outskirts have lower metallicity than their inner regions (e.g. Pilkington et al., 2012; Ju et al., 2025). However, a similar argument could be made for e.g. core-collapse gamma-ray bursts, which are typically not observed at large offsets. Regardless of offsets, the host metallicities of LFBOTs extend above the sharp cut-off predicted for very massive star core-collapse events (see Sec 3). Furthermore, metallicities from SED fitting should be treated with caution (e.g. Nersesian et al., 2025). For example, the mass-metallicity relation for the host of AT 2023fhn suggests a metallicity of 0.5 Z⊙, versus the fitted value of 0.1 Z⊙. In any case, the metallicity distributions as presented in Fig. 2 are inconsistent, but the possibility of low-Z pockets on spatial scales smaller than probed by current measurements has the potential to resolve this discrepancy.
@@ 5.3 A potential source of r-process nucleosynthesis
Massive black holes and their accretion discs, such as those discussed here, have also been predicted as possible sites of ‘super-kilonovae’. In this scenario, the density in a high accretion disc becomes high enough to trigger r-process nucleosynthesis (Siegel et al., 2019, 2022; Dean & Fernández, 2024; Rastinejad et al., 2024; Agarwal et al., 2025). The possibility of such a channel has implications for where and when r-process elements are produced in the Universe (Nugent et al., 2025). Given the black hole and accreted masses involved in the failed supernova interpretation of LFBOTs, they may be promising sites for r-process production. The volumetric rate of LFBOTs in the redshift range considered in this work (Ho et al., 2023b, 0.1% of the CCSN rate, or 70 yr-1 Gpc-3) is compatible with the rate of super-kilonovae predicted by Siegel et al. (2022) of 10–100 yr-1 Gpc-3. We have shown that the failed-supernova interpretation of LFBOTs naturally favours events forming black holes greater than M⊙, while super-KNe have been predicted in events producing black holes greater than M⊙ (Siegel et al., 2022). If the true rate of LFBOTs is just a factor of few lower than the current 70 yr-1 Gpc-3 best estimate, then the BH mass cut required to match the LFBOT rate moves up to the pair instability mass gap (see Figure 1), and we have a 1:1 mapping between LFBOT and super-KN progenitors. In any case, there is surely substantial overlap in the populations. Furthermore, a near-infrared excess - a hallmark of radiative reprocessing by r-process elements - was observed on a timescale of weeks following both AT 2018cow (Perley et al., 2019) and AT2024wpp (Pursiainen et al., 2025). This has been described either as a dust echo or due to free-free opacity effects outside the optical photosphere in both events (Metzger & Perley, 2023; LeBaron et al., 2025), but no spectroscopic characterisation of this feature in an LFBOT has yet been obtained. Future (James Webb Space Telescope) near-infrared spectroscopic observations can distinguish between an r-process element-forming super-kilonova and a featureless dust echo or free-free emission.
@@ 5.4 Further implications
There are several other observations associated with LFBOTs which any model invoked to explain them must satisfy, assuming that LFBOTs are homogenous class. For example, extremely high velocity (), blue-shifted hydrogen lines were observed in the CSS161010 (Gutiérrez et al., 2024), perhaps most easily explained as out-flowing streams of material in a tidal disruption event, although these have only been observed in CSS161010 thus far. Polarimetric observations have found that LFBOTs start highly polarised and trend towards low polarisation on a timescale of days (Maund et al., 2023; Pursiainen et al., 2025). This can plausibly be explained in a failed supernova context by (i) an accretion disc producing the high early-time polarisation and (ii) a rapid transition to an spherical outflow (Pursiainen et al., 2025), perhaps caused by outflows from highly super-Eddington accretion (e.g. Coppejans et al., 2020). The origin of the giant optical flares observed from AT2022tsd in the months after the event is unclear, but timescale arguments constrain the scale of the emitting region to the outer regions of an accretion disc, for a stellar to intermediate mass black hole central engine (Ho et al., 2023a).
@@ 6 Conclusions
We show that the mass loss rates and stellar wind speeds associated with these stellar models correspond to dense circumstellar media, consistent with the lower end of LFBOT radio observations, but struggling to reproduce the most dense environments inferred. However, as bpass only runs until the end of core carbon burning, possible large mass loss phases occurring after carbon burning could explain the high-densities,
@@
The long-lived UV emission seen in AT 2018cow can be reproduced in the failed supernova interpretation given black hole masses M⊙ and several solar masses of fallback accretion.
@@
We therefore conclude that the failed supernova scenario is a plausible explanation for at least some LFBOTs, but it hinges on the measured metallicities at the LFBOT site and circumstellar densities. The true feasibility of this progenitor channel clearly depends on the detailed physics of very massive star core-collapse, late-stage mass loss, and the amount of mass which falls back or is ejected.

### v2

@@ 4.1 H and He-poor spectra
LFBOTs display H and He poor spectra at early times (e.g. 2018ApJ...865L...3P). The progenitor star, or star being tidally disrupted, therefore must have a He or He-poor composition. In the TDE scenario, a white dwarf is favoured (2019MNRAS.487.2505K). For a BH-stellar merger, the stellar object must be a Wolf-Rayet star (2022ApJ...932...84M). In the scenario explored in this paper, the median H and He mass fractions in the ‘leftover mass’222I.e. the stellar mass leftover after remnant formation. In a successful supernova with no fallback, this would be the ejecta mass. of the models selected as progenitors of BHs with M⊙ are X0.003 and Y0.08 respectively (where the uncertainties are defined by the 16th and 84th percentiles). The distributions of Xsurf and Ysurf are shown in Figure 3. 69% of our selected models are classified as Wolf-Rayet stars given the Xsurf values and high surface temperatures of log10(T/K)4.45 (2017PASA...34...58E). The outer layers of these stars therefore meet the criteria for being H and He poor, and successfully exploding stars with such low X and Y surface mass fractions are expected to produce type Ic supernovae (i.e. SNe without detectable H and He lines, e.g. 2012MNRAS.424.2139D; 2013MNRAS.436..774E; 2020MNRAS.491.3479C).
@@ 4.2 Dense circumstellar media and radio emission
A key characteristic of LFBOTs is their slow-rising, luminous radio emission, attributed to self-absorbed synchrotron emission from a blast-wave expanding through the circumstellar medium (CSM). These blast-waves are mildly relativistic (e.g. 2020ApJ...895L..23C), and probe the CSM at radii of 1016–1017 cm. LFBOT radio emission is well described by blast-wave propagation through dense CSM profiles defined by values of 1–100, in units of M⊙ yr-1/1000 km s-1. For Wolf-Rayet wind speeds of order 1000 km s-1, radii of 1016–1017 cm corresponds to mass loss in the final decades before explosion. We calculate the mean final values for the selected models, noting that bpass only runs until the end of core carbon burning. Mass loss rates are provided as a bpass output. We adopt standard bpass wind speed prescriptions, namely those of 2001A&A...369..574V, and 2000A&A...360..227N for Wolf-Rayets (we refer the reader to 2017PASA...34...58E; 2022MNRAS.515.2591C, for more details). We find a median final = (0.04) M⊙ yr-1/1000 km s-1. The full distribution of is shown in Figure 5. The bulk of our progenitor mass-loss rate and wind speed values lie below the values inferred from LFBOT radio observations, which span the range 1-100 in these units (e.g. 2022ApJ...932..116H; 2022ApJ...926..112B, under the assumption of and ). We note that mass loss in the final decades before core-collapse may be significant and eruptive, plausibly explaining the non- profiles observed (e.g. 2014ARA&A..52..487S). However, because the stellar evolution models used end after core carbon burning, a full modelling of the CSM around these stars on the relevant scales is beyond the scope of this paper.
@@ 4.3 Long-lived ultraviolet plateaus
Finally, we ask whether the selected stellar models can produce a long-lived tail of UV emission lasting at least several years after the event, as observed in AT2108cow (2022MNRAS.512L..66S; 2023MNRAS.519.3785S; 2023ApJ...955...43C; 2023MNRAS.525.4042I; 2025MNRAS.544L.108I). This has been interpreted as due to a long-lived accretion disc, reminiscent of the emission seen in TDEs (2022ApJ...932...84M; 2023MNRAS.525.4042I; 2025MNRAS.541..429M). To see whether our models can produce AT 2018cow-like UV plateaus through this mechanism, we first examine the remaining mass in the selected bpass models which survives outside the nascent BHs (i.e., the pre-core-collapse mass minus the initial BH mass). In Figure 4 we show 2D histograms of these masses at each metallicity, for the fiducial 2006ApJ...638L..63L CSFH. For more details on the BH mass distribution in bpass, we refer the reader to 2017PASA...34...58E, 2022MNRAS.511.1201G and 2023MNRAS.520.5724B. More detail on the progenitors, in terms of temperature, luminosity and radius at the end of the models (i.e. the end of core carbon burning) is provided in Section A of the Appendix.
@@
As outlined in Section 2, bpass calculates BH masses by injecting a standard supernova energy of 1051 erg to determine the mass of the material which is unbound. The remainder is the remnant mass. In a failed supernova, the ‘leftover’ material is not successfully ejected. It may therefore be reasonable to assume that all of the leftover material falls back and is accreted instead, in line with other prescriptions for final BH masses in the massive to very massive star regime, which have close to a one-to-one relation between the pre-core-collapse mass and final remnant mass (after fallback) for stripped envelope progenitors (2012ApJ...749...91F; 2023MNRAS.520.5724B). With the BH mass MBH, leftover mass Mleftover, total pre-collapse stellar mass M⋆, radius R⋆ and rotational velocity v⋆ as inputs, we determine an initial disc radius as follows,
where and are the fraction of the pre-collapse angular momentum and leftover mass which go into the disc, respectively. A derivation of Equation 1 is provided in Appendix B. Since stellar rotation is not explicitly tracked in bpass, we have assumed the pre-collapse rotational velocity to be 10% of the breakup (critical) velocity (massive stars in low metallicity environments such as the Magellanic Clouds are observed to be rotating at 10–20% of their critical velocity e.g. 2013A&A...560A..29R; 2015A&A...580A..92R). We proceed to use the model of 2020MNRAS.492.5655M; 2025MNRAS.544.2225M to calculate UV light-curves for every selected stellar model. The final parameter required (beyond the disc mass, black hole mass and radial scale) is the initial evolutionary timescale of the disc. We follow standard SS73 theory and take , where is a dimensionless factor, and is the disc aspect ratio. Anticipating a thick disc (for this likely initially super-Eddington flow) we sample in the range , inducing scatter in our simulated light curves. This fixes all parameters in the disc model.
We adopt two example scenarios, and show the predicted UV lightcurves in Figure 6 for the F225W Wide Field Camera 3 (WFC3) Hubble Space Telescope filter. In first two, only 10% of the available angular momentum and mass is transferred to the disc. In the second extreme case, all of the angular momentum and mass goes into the disc. We acknowledge that the details details of supernova fallback, such as its asphericity and the fraction which falls back versus being ejected, are not captured in the models or our simple parametrisation. This motivated our choice of two scenarios which nevertheless represent a wide range of possible fallback masses and angular momenta, including the hypothetical maximum case. The total leftover mass is also uncertain (and depends on the injected energy at core-collapse), for a discussion of the impact of this we refer the reader to Section 5.3.
The resulting light-curves can only match the observed late-time F225W flux for AT 2018cow if a large fraction of the leftover mass and angular momentum are transferred to the disc (see also Omand et al. in prep). The initial disc radii predicted from these models are considerably smaller than the 30-40 R⊙ inferred from the late-time fits to AT2018cow (2023MNRAS.525.4042I; 2024ApJ...963L..24M). This is because the majority of the progenitors are compact stars (e.g. Wolf-Rayets, see Figure 7 of Appendix A). However, the discs spread rapidly (as t2/3), reaching 10 R⊙ within a day and 100 R⊙ within 1000 days. We note that the factors facc and fj,disc cancel, such that the two cases adopted return the same initial disc radii. The growth of the outer disc radius will be less pronounced if angular momentum is transported away from the disc by winds (potentially likely at early times), see Appendix B. We find that covers a wide range of 0.5-106 days. Only for the shortest of these timescales can the same accretion mechanism plausibly explain the rapid early-time rise in LFBOT lightcurves. Otherwise, we acknowledge that some other process must drive the initial rise and peak.
@@
The failed SN interpretation of LFBOTs is predicated on the notion that some non-negligible fraction of the total stellar mass does not participate in the prompt BH formation and instead falls back onto the nascent BH. If the core-collapse of such massive stars actually results in the entire stellar mass being promptly engulfed by the event horizon, there will no material to accrete, ruling out failed supernovae as a viable channel for LFBOTs. However, in our fiducial models this is not the case (see Figure 4, as well as a discussion of the explodability assumptions in Section 5.3).
@@ 5.1 LFBOT environments
In the failed SN scenario, we would expect LFBOTs to be strongly associated with star-forming regions, as the short-lived massive progenitors are not expected to travel far from their birth sites. Although the LFBOT sample is small, it appears as though LFBOTs occupy diverse environments within their hosts, with at least two events at moderate to large host-normalised offsets (2024MNRAS.527L..47C; 2025MNRAS.537.3298P). The apparent preference for LFBOTs to occur in low metallicity environments may help in this regard: in general, galactic outskirts have lower metallicity than their inner regions (e.g. 2012A&A...540A..56P; 2025ApJ...978L..39J). However, a similar argument could be made for e.g. core-collapse gamma-ray bursts, which are typically not observed at large offsets. Regardless of offsets, the host metallicities of LFBOTs extend above the sharp cut-off predicted for very massive star core-collapse events (see Sec 3). Furthermore, metallicities from SED fitting should be treated with caution (e.g. 2025A&A...695A..86N). For example, the mass-metallicity relation for the host of AT 2023fhn suggests a metallicity of 0.5 Z⊙, versus the fitted value of 0.1 Z⊙. In any case, the metallicity distributions as presented in Fig. 2 are inconsistent, but the possibility of low-Z pockets on spatial scales smaller than probed by current measurements has the potential to resolve this discrepancy.
@@ 5.2 Initial mass function
5.3 Explodability and remnant mass assumptions
The results in this paper depend on the BH masses formed and the mass available in the ‘leftover’ material for fallback. The models adopted calculated BH masses by injecting 1051 erg of energy and determining the mass which which remains bound. If the actual energy injection in failed events is 1051 erg (as expected, e.g. 2018MNRAS.476.2366F; 2023MNRAS.526..152K) then the BH masses will be greater. However, the ejecta/leftover masses are typically not more than 10% of the BH mass, so the most they could increase by is 10, which would not significantly change our conclusions. Ultimately, the models predict similar final remnant masses to other widely used prescriptions in the high mass regime (e.g. 2012ApJ...749...91F, see also Figure 8 of Appendix A).
There is also the question of which stars explode and which undergo a failed collapse. Our assumption that all BH-forming events fail is an upper limit on the rate of failed SNe. The predicted rate of all failed SNe under this assumption is 0.2 of the CCSN (NS) formation rate, which is compatible with upper limits on the rate of failed SNe from searches for faint, long-lived red transients (thought to be a likely outcome for failed SNe in the absence of any additional energy injection from a disc, 2022MNRAS.514.1188B).
However, we are likely overestimating the failed SN rate, given recent hydrodynamical simulations and suggestions that there may not be a missing red supergiant problem (e.g. 2025ApJ...987..164B; 2025ApJ...979..117B). If failed SNe only occur in the 12-15 M⊙ (ZAMS mass) window where failed events are reliably predicted to occur in these simulation (and assuming NS formation everywhere outside this range unless a PISN occurs), then we would expect a ratio of failed events to CCSNe of 0.15. The lower masses (higher up the IMF) counteract the narrower mass range, producing a similar ratio to our fiducial 30–40 M⊙ BH mass cut. Such a scenario could therefore reproduce the LFBOT rate but not the inferred BH masses. The prospect of incorporating more physically motivated prescriptions for remnant and ejecta mass distributions in the framework of population synthesis (e.g. 2020MNRAS.499.2803P; 2025A&A...700A..20M) could help refine results such as these in future.
@@ 5.4 A potential source of r-process nucleosynthesis
Massive black holes and their accretion discs, such as those discussed here, have also been predicted as possible sites of ‘super-kilonovae’. In this scenario, the density in a high accretion disc becomes high enough to trigger r-process nucleosynthesis (2019Natur.569..241S; 2022ApJ...941..100S; 2024PhRvD.110h3024D; 2024ApJ...968...14R; 2025arXiv250315729A). The possibility of such a channel has implications for where and when r-process elements are produced in the Universe (2025ApJ...982..144N). Given the black hole and accreted masses involved in the failed supernova interpretation of LFBOTs, they may be promising sites for r-process production. The volumetric rate of LFBOTs in the redshift range considered in this work (2023ApJ...949..120H, 0.1% of the CCSN rate, or 70 yr-1 Gpc-3) is compatible with the rate of super-kilonovae predicted by 2022ApJ...941..100S of 10–100 yr-1 Gpc-3. We have shown that the failed-supernova interpretation of LFBOTs naturally favours events forming black holes greater than M⊙, while super-KNe have been predicted in events producing black holes greater than M⊙ (2022ApJ...941..100S). If the true rate of LFBOTs is just a factor of few lower than the current 70 yr-1 Gpc-3 best estimate, then the BH mass cut required to match the LFBOT rate moves up to the pair instability mass gap (see Figure 1), and we have a 1:1 mapping between LFBOT and super-KN progenitors. In any case, there is surely substantial overlap in the populations. Furthermore, a near-infrared excess - a hallmark of radiative reprocessing by r-process elements - was observed on a timescale of weeks following both AT 2018cow (2019MNRAS.484.1031P) and AT2024wpp (2025MNRAS.537.3298P). This has been described either as a dust echo or due to free-free opacity effects outside the optical photosphere in both events (2023ApJ...944...74M; 2025ApJ...993L...6N), but no spectroscopic characterisation of this feature in an LFBOT has yet been obtained. Future (James Webb Space Telescope) near-infrared spectroscopic observations could potentially distinguish between an r-process element-forming super-kilonova and a featureless dust echo or free-free emission. However, in order for a UV-emitting disc and IR-emitting super-kilonova to be simultaneously observable, an anisotropic geometry which allows simultaneous viewing of both photospheres would be required.
@@ 5.5 Further implications
There are several other observations associated with LFBOTs which any model invoked to explain them must satisfy, assuming that LFBOTs are homogenous class. For example, extremely high velocity (), blue-shifted hydrogen lines were observed in the CSS161010 (2024ApJ...977..162G), perhaps most easily explained as out-flowing streams of material in a tidal disruption event, although these have only been observed in CSS161010 thus far. Polarimetric observations have found that LFBOTs start highly polarised and trend towards low polarisation on a timescale of days (2023MNRAS.521.3323M; 2025MNRAS.537.3298P). This can plausibly be explained in a failed supernova context by (i) an accretion disc producing the high early-time polarisation and (ii) a rapid transition to an spherical outflow (2025MNRAS.537.3298P), perhaps caused by outflows from highly super-Eddington accretion (e.g. 2020ApJ...895L..23C). The origin of the giant optical flares observed from AT2022tsd in the months after the event is unclear, but timescale arguments constrain the scale of the emitting region to the outer regions of an accretion disc, for a stellar to intermediate mass black hole central engine (2023Natur.623..927H).
@@ 6 Conclusions
We show that the mass loss rates and stellar wind speeds associated with these stellar models correspond to dense circumstellar media, but which nevertheless only overlap with the lower end of LFBOT radio observations, and struggle to reproduce the typical circumstellar densities inferred. However, as bpass only runs until the end of core carbon burning, possible large mass loss phases occurring after carbon burning could explain the high-densities,
@@
The long-lived UV emission seen in AT 2018cow can be reproduced in the failed supernova interpretation given black hole masses M⊙, but this requires (i) several solar masses of fallback accretion and (ii) a high fraction of the pre-collapse stellar angular momentum to be transferred to the disc,
@@
We therefore conclude that the failed supernova scenario is a plausible explanation for at least some LFBOTs, but it hinges on the measured metallicities at the LFBOT site and the origin of the dense circumstellar media. The true feasibility of this progenitor channel clearly depends on the detailed physics of very massive star core-collapse, late-stage mass loss, and the amount of mass which falls back or is ejected.